# Comparison of Structure-Preserving Methods for the Cahn-Hilliard-Navier-Stokes Equations

### v1

@@ 1 Introduction
While Theorem 1.1, proven in (Elliott and Garcke, 1996, Thm. 1), guarantees a weak maximum principle at the continuum level for a degenerate mobility, achieving discrete boundedness remains a significant challenge. The standard Finite Element Method (FEM), symmetric weighted interior penalty (SWIP), and symmetric interior penalty Galerkin (SIPG) discretizations fail to preserve boundedness without additional stabilization Gunnarsson and Klöfkorn (2026), even when using the degenerate mobility $M$. Recent work, however, has demonstrated that boundedness can be numerically achieved through carefully designed limited Galerkin schemes, including FEM-L for FEM, and for DG SIPG-L, and SWIP-L Gunnarsson and Klöfkorn (2026) with degenerate CH.
In this work, we extend these structure-preserving methods by introducing new mobility fluxes using the Discontinuous Galerkin (DG) formulations with intersection-wise mobility treatments. These extend the SWIP-L and SIPG-L methods introduced in Gunnarsson and Klöfkorn (2026), by allowing for better control of the coercivity-stability balance.
@@
where $\rho(\psi):=\frac{1+\psi}{2}\rho_{1}+\frac{1-\psi}{2}\rho_{2}$ is the density, $\boldsymbol{J}:=\frac{\rho_{2}-\rho_{1}}{2}Pe^{-1}M(\psi)\nabla\upsilon$ is the mass flux, $P$ is the pressure, $\mu(\psi):=\frac{1+\psi}{2}\mu_{1}+\frac{1-\psi}{2}\mu_{2}$ is the viscosity, and $D(\boldsymbol{u})=\frac{1}{2}(\nabla\boldsymbol{u}+\nabla\boldsymbol{u}^{T})$ is the rate of deformation tensor, where $\mathbb{I}$ denotes the identity tensor. Moreover, $\rho_{j}$ and $\mu_{j}$ are the characteristic densities and viscosities for phases $j=1,2$, respectively. We equip the boundary conditions $\mathbf{n}\cdot\boldsymbol{u}|_{\partial\Omega}=\mathbf{n}\cdot\nabla P|_{\partial\Omega}=0$. We note that the expected bounds for the physical mass density $\rho\in[\rho_{1},\rho_{2}]$ follows directly from boundedness of the phase-field $\psi\in[-1,1]$.
@@ 2 Discretization
where $\mathbb{P}^{p}({\mathcal{T}_{h}})$ denotes a polynomial space of order at most $p$ on the grid $\mathcal{T}_{h}$}. We also permit $p$-adaptivity over the mesh $\mathcal{T}_{h}$}, and as such, the polynomial order $p$ may vary across elements $K$, and we denote the local polynomial order as $p_{K}$. We denote the maximal permitted polynomial order as $p_{\max}$, and the minimal order as $p_{\min}$. We describe the application further in Sec. 2.2.
@@ 2.1 Discontinuous Galerkin
Consider an intersection $e\in\Gamma_{i}$ shared by two elements $K^{-}_{e}$ and $K^{+}_{e}$. Then for $K\in{\mathcal{T}_{h}}$ there exists a constant $C_{K}>0$ independent of the mesh width $h$ such that for all $\varphi_{h}\in V^{p}_{h}$ for $p>0$ and $e\in\Gamma_{i}$ the following inequality holds:
@@
where, in the special case of $K_{e}$ being a quadrilateral or triangle, $C_{K_{e}^{\pm}}=\frac{p_{K}(p_{K}+d-1)}{d}$ where $p_{K}$ is the local polynomial degree (see further treatments and formulations in Ainsworth and Rankin (2009) and Epshteyn and Rivière (2007)).
@@
Under the same setting as in Lemma 1 and for $\varphi_{h}\in V_{h}^{p}$ the following inequality holds for all $e\in\Gamma_{i}$:
@@
where for the special case when $K_{e}$ is a quadrilateral or triangle:
@@
For the Navier-Stokes coupling we follow exactly the same formulation as in Gunnarsson and Klöfkorn (2026) with a post-processing step. Our modification includes projecting the FEM velocity field $\boldsymbol{u}$ to a Brezzi-Douglas-Marini space of first order to obtain a divergence-free velocity field $\tilde{\boldsymbol{u}}$.
@@ 3 Numerics
In this section, we present numerical results for the proposed schemes for both the standalone CH equations and the coupled CH-NS system. We first consider Ex. 1 to verify the optimal convergence rates in 2D and 3D. We then investigate the performance of the proposed schemes for merging droplets in Ex. 2, and finally consider the rotating droplet configuration in Ex. 3 for the coupled CH-NS system to demonstrate the structure-preserving properties and $hp$-adaptivity. The numerical simulations are implemented in Dune-Fem Bastian et al. (2021); Dedner and Klöfkorn (2021) at version $2.11$ with its Python interface Dedner et al. (2020).
@@
Consider the domain $\Omega=[0,1]^{d}$ and the following artificial solution to the CH equations $\psi(\boldsymbol{x},0)=A\prod_{i=1}^{d}\cos(4\pi\boldsymbol{x}_{i})$ where $A\in(0,1]$ is an amplitude parameter and $\boldsymbol{x}_{i}$ is the $i$-th component of the spatial coordinate $\boldsymbol{x}$. We then consider the following forcing term $S(\boldsymbol{x},t)=Pe^{-1}\nabla\cdot(M(\psi)(Cn^{2}\Delta\psi-W^{\prime}(\psi)))$ added to the right-hand side of Eq.(22), where $Cn=0.1$, $Pe=0.3$ are the Cahn and Peclet numbers respectively. The simulation is run for $t\leq T=0.01$.
@@
In Table 1, we present the $L^{2}$ and $H^{1}$ errors and EOC for Ex. 1 for the different schemes with amplitudes $A=0.1$ and $A=0.6$ and $\beta=5$. The results are consistent with the theoretically expected convergence rates. Additionally, SWIPD-L/SIPGD-L and SWIP-L/SIPG-L yield identical errors and EOC, which is consistent with the fact that the mobility flux does not affect the consistency of the scheme, but only the stability and coercivity. Similar behavior for SIPG-L and SWIP-L was also observed in Gunnarsson and Klöfkorn (2026). We also tested if the same convergence rates hold in 3D, and the results are presented in Table 2 for $A=0.1$ and $\tau=10^{-3}\cdot 2^{-\frac{N}{16}}$. The observed rates confirm the theoretical predictions.
@@
For Exs. 2 and 3, we consider an $hp$-adaptive formulation with $h_{\min}=\frac{1}{256}$ and $h_{\max}=\frac{1}{128}$ using the SWIPD-L scheme and a prescribed maximum polynomial degree of $p_{max}$ and $p_{min}\geq 0$. For the time increments, following Gunnarsson and Klöfkorn (2026), we set $\tau=10^{-3}$ for Ex. 2 and $\tau=5\cdot 10^{-4}$ for Ex. 3. We simulate with different order levels of $p$ for $p$-adaptivity and also compare to only the $h$-adaptive formulation, i.e., $p=p_{max}$ for all elements $K\in{\mathcal{T}_{h}}$. We observe that convergence was not achieved for Ex. 2 with $p=2$ for the $h$-adaptive formulation; convergence was only obtained, in that case, when $\psi_{h}||_{L^{\infty}(K)}\lessapprox 1$, as was also stressed in Gunnarsson and Klöfkorn (2026) regarding careful selection of the initial data. To the best of the authors’ knowledge there is no clear proof or indication on how to find a safe initial value.
@@
Figure 3 shows the energy, mass, and boundedness for Ex. 2 using the $hp$-adaptive formulation of SWIPD-L with for various $p$-levels for $hp$-adaptivity. The energy is monotonically non-increasing, mass is conserved, and the phase-field remains bounded, consistent with the theoretical properties. The results are similar to those obtained using only $h$-adaptivity with $p=1$, and for $p_{\max}=1$ with $hp$-adaptivity, which suggests that the $hp$-adaptive formulation preserves accuracy alongside the $h$-adaptive formulation, while providing a significant reduction in complexity and degrees of freedom.
@@
Figure 4 shows that the energy is non-increasing, mass is conserved, and the phase-field remains bounded for Ex. 3 using the $hp$-adaptive formulation of SWIPD-L with and without $p$-adaptivity for the studied levels. The results are similar to those obtained using only $h$-adaptivity for the studied $p$-levels, which further indicates that the $hp$-adaptive formulation preserves physical accuracy alongside to the $h$-adaptive formulation, while providing a significant reduction in complexity and degrees of freedom, similarly to our discussion for Ex. 2. We also note that for $p=2$ the mass difference in Fig. 4 is slowly increasing due to the nonlinear tolerance $\epsilon$ (for these simulations we picked $\epsilon=5\cdot 10^{-16}$) having to be picked slightly larger to avoid solver stagnation. These minor mass drifts are a known issue with potential deviations scaling as the non-linear tolerance as noted in Gunnarsson and Klöfkorn (2026). Regardless, as is seen in Figs. 3 and 4 showing that the mass change is negligible in comparasion to the lines at $\Delta m=\pm\frac{T}{\tau}\epsilon$ for $\epsilon=2\cdot 10^{-16}$ for Ex. 2 and $\epsilon=5\cdot 10^{-16}$ for Ex. 3.
@@ 4 Summary and Outlook
We have proposed a new class of DG schemes for the CH equations with degenerate mobility has been proposed, which are provably coercive and stable under a certain condition on the mobility flux for the penalty term. The introduced mobility flux formulations for the penalty term include SIPGD-L, which employs the intersection maximum of the mobility, and SWIPD-L, which uses a harmonic average. These mobility fluxes yield similar estimates for the coercivity condition and accuracy as previous schemes. Numerical experiments illustrate the performance of the proposed schemes, demonstrating that they achieve optimal convergence rates and that the harmonic mobility flux provides potentially significant improvements in stability and accuracy compared to the arithmetic mean. An $hp$-adaptive formulation has also been presented, showing significant reductions in complexity and degrees of freedom compared to the $h$-adaptive formulation while maintaining similar accuracy and physical fidelity.
Future work includes extending the theoretical analysis to provide rigorous proofs of boundedness for the SWIPD-L scheme in the broken Sobolev (DG) setting, exploiting the connection between the harmonic average and the minimum to extend Theorem 1.1. Additionally, more challenging test cases, such as the rising bubble benchmark, will further validate the performance of the proposed schemes for coupled CHNS systems and assess adaptivity effects under more demanding conditions.

### v2

@@ 1 Introduction
While Theorem 1.1, proven in (Elliott and Garcke, 1996, Thm. 1), guarantees a weak maximum principle at the continuum level for a degenerate mobility, achieving discrete boundedness remains a significant challenge. The standard Finite Element Method (FEM), symmetric weighted interior penalty (SWIP), and symmetric interior penalty Galerkin (SIPG) discretizations fail to preserve boundedness without additional stabilization Gunnarsson and Klöfkorn (2026), even when using the degenerate mobility $M$. Recent work, however, has demonstrated that boundedness can be numerically achieved through carefully designed limited Galerkin schemes, including FEM-L for FEM, and for Discontinuous Galerkin (DG) schemes with SIPG-L, and SWIP-L Gunnarsson and Klöfkorn (2026) with the degenerate CH equations.
In this work, we extend these structure-preserving methods by introducing new mobility fluxes using the DG formulations with intersection-wise mobility treatments. These extend the SWIP-L and SIPG-L methods introduced in Gunnarsson and Klöfkorn (2026), by allowing for better control of the coercivity-stability balance.
@@
where $\rho(\psi):=\frac{1+\psi}{2}\rho_{1}+\frac{1-\psi}{2}\rho_{2}$ is the density, $\boldsymbol{J}:=\frac{\rho_{2}-\rho_{1}}{2}Pe^{-1}M(\psi)\nabla\upsilon$ is the mass flux, $P$ is the pressure, $\mu(\psi):=\frac{1+\psi}{2}\mu_{1}+\frac{1-\psi}{2}\mu_{2}$ is the viscosity, and $D(\boldsymbol{u})=\frac{1}{2}(\nabla\boldsymbol{u}+\nabla\boldsymbol{u}^{T})$ is the rate of deformation tensor, where $\mathbb{I}$ denotes the identity tensor. Moreover, $\rho_{j}$ and $\mu_{j}$ are the characteristic densities and viscosities for phases $j=1,2$, respectively. We equip the boundary conditions $\mathbf{n}\cdot\boldsymbol{u}|_{\partial\Omega}=\mathbf{n}\cdot\nabla P|_{\partial\Omega}=0$. We note that the expected bounds for the physical mass density $\rho\in[\rho_{1},\rho_{2}]$ follow directly from boundedness of the phase-field $\psi\in[-1,1]$.
@@ 2 Discretization
where $\mathbb{P}^{p}(K)$ denotes a polynomial space of order at most $p$ on an element $K$. We also permit $p$-adaptivity over the mesh $\mathcal{T}_{h}$}, and as such, the polynomial order $p$ may vary across elements $K$, and we denote the local polynomial order as $p_{K}$. We denote the maximal permitted polynomial order as $p_{\max}$, and the minimal order as $p_{\min}$. We describe the application further in Sec. 2.2.
@@ 2.1 Discontinuous Galerkin
Consider an intersection $e\in\Gamma_{i}$ shared by two elements $K^{-}_{e}$ and $K^{+}_{e}$. Then there exists a constant $C_{K_{e}^{\pm}}>0$ independent of the mesh width $h$ such that for all $\varphi_{h}\in V^{p}_{h}$ for $p>0$ and $e\in\Gamma_{i}$ the following inequality holds:
@@
where, in the special case of $K_{e}$ being a quadrilateral or triangle, $C_{K_{e}^{\pm}}=\frac{p(p+d-1)}{d}$ (see further treatments and formulations in Ainsworth and Rankin (2009) and Epshteyn and Rivière (2007)).
@@
Under the same setting as in Lemma 1 and for $\varphi_{h}\in V_{h}^{p_{\max}}$ for ${p_{\max}}>0$. the following inequality holds for all $e\in\Gamma_{i}$:
@@
where for the special case where $K^{\pm}_{e}$ are quadrilaterals or triangles:
@@
For the NS coupling we follow exactly the same formulation as in Gunnarsson and Klöfkorn (2026) with a post-processing step. Our modification includes projecting the FEM velocity field $\boldsymbol{u}$ to a Brezzi-Douglas-Marini space of first order to obtain a divergence-free velocity field $\tilde{\boldsymbol{u}}$.
@@ 3 Numerics
In this section, we present numerical results for the proposed schemes for both the standalone CH equations and the coupled CH-NS system. We first consider Ex. 1 to verify the optimal convergence rates in 2D and 3D. We then investigate the performance of the proposed schemes for merging droplets in Ex. 2, and finally consider the rotating droplet configuration in Ex. 3 for the coupled CH-NS system to demonstrate the structure-preserving properties and $hp$-adaptivity. The numerical simulations are implemented in Dune-Fem Bastian and et al. (2021); Dedner and Klöfkorn (2021) at version $2.11$ with its Python interface Dedner et al. (2020).
@@
Consider the domain $\Omega=[0,1]^{d}$ and the following artificial solution to the CH equations $\psi(\boldsymbol{x},0)=A\prod_{i=1}^{d}\cos(4\pi\boldsymbol{x}_{i})$ where $A\in(0,1]$ is an amplitude parameter and $\boldsymbol{x}_{i}$ is the $i$-th component of the spatial coordinate $\boldsymbol{x}$. We then consider the following forcing term $S(\boldsymbol{x},t)=Pe^{-1}\nabla\cdot(M(\psi)\nabla(W^{\prime}(\psi)-Cn^{2}\Delta\psi))$ added to the right-hand side of Eq. (22), where $Cn=0.1$, $Pe=0.3$ are the Cahn and Peclet numbers respectively. The simulation is run for $t\leq T=0.01$.
@@
In Table 1, we present the $L^{2}$ and $H^{1}$ errors and EOC for Ex. 1 for the different schemes with amplitudes $A=0.1$ and $A=0.6$ given $\beta=5$. The results are consistent with the theoretically expected convergence rates. Additionally, SWIPD-L/SIPGD-L and SWIP-L/SIPG-L yield identical errors and EOC, which is consistent with the fact that the mobility flux does not affect the consistency of the scheme, but only the stability and coercivity. Similar behavior for SIPG-L and SWIP-L was also observed in Gunnarsson and Klöfkorn (2026). We also tested if the same convergence rates hold in 3D, and the results are presented in Table 2. The observed rates confirm the theoretical predictions.
@@
For Exs. 2 and 3, we consider an $hp$-adaptive formulation with $h_{\min}=\frac{1}{256}$ and $h_{\max}=\frac{1}{128}$ using the SWIPD-L scheme and a prescribed maximum polynomial degree of $p_{\max}$} and ${p_{\min}}\geq 0$. For the time increments, following Gunnarsson and Klöfkorn (2026), we set $\tau=10^{-3}$ for Ex. 2 and $\tau=5\cdot 10^{-4}$ for Ex. 3. We simulate with different order levels of $p$ for $p$-adaptivity and also compare to only the $h$-adaptive formulation, i.e., $p={p_{\max}}$ for all elements $K\in{\mathcal{T}_{h}}$. We observe that convergence was not achieved for Ex. 2 with $p=2$ for the $h$-adaptive formulation; convergence was only obtained, in that case, when $\psi_{h}||_{L^{\infty}(K)}\lessapprox 1$, as was also stressed in Gunnarsson and Klöfkorn (2026) regarding careful selection of the initial data. To the best of the authors’ knowledge there is no clear proof or indication on how to find a safe initial value.
@@
Figure 3 shows the energy, mass, and boundedness for Ex. 2 using the $hp$-adaptive formulation of SWIPD-L for various $p$-levels with $hp$-adaptivity. The energy is monotonically non-increasing, mass is conserved, and the phase-field remains bounded, consistent with the theoretical properties. The results are similar to those obtained using only $h$-adaptivity with $p=1$, and for $p_{\max}=1$ with $hp$-adaptivity, which suggests that the $hp$-adaptive formulation preserves accuracy alongside the $h$-adaptive formulation, while providing a significant reduction in complexity and degrees of freedom.
@@
Figure 4 shows that the energy is non-increasing, mass is conserved, and the phase-field remains bounded for Ex. 3 using the $hp$-adaptive formulation of SWIPD-L both with and without $p$-adaptivity. The results are similar to those obtained using only $h$-adaptivity for the studied $p$-levels, which further indicates that the $hp$-adaptive formulation preserves physical accuracy alongside the $h$-adaptive formulation, while providing a significant reduction in complexity and degrees of freedom, similarly to our discussion for Ex. 2. We also note that for $p=2$ the mass difference in Fig. 4 is slowly increasing due to the nonlinear tolerance $\epsilon$ (for these simulations we picked $\epsilon=5\cdot 10^{-16}$) having to be picked slightly larger to avoid solver stagnation. These minor mass drifts are a known issue with potential deviations scaling as the non-linear tolerance as noted in Gunnarsson and Klöfkorn (2026). Regardless, as shown in Figs. 3 and 4, the mass change is negligible compared to the lines at $\Delta m=\pm\frac{T}{\tau}\epsilon$ for $\epsilon=2\cdot 10^{-16}$ for Ex. 2 and $\epsilon=5\cdot 10^{-16}$ for Ex. 3.
@@ 4 Summary and Outlook
A new class of DG schemes for the CH equations with degenerate mobility has been proposed, which is provably coercive and stable under a certain condition on the mobility flux for the penalty term. The introduced mobility flux formulations for the penalty term include SIPGD-L, which employs the intersection maximum of the mobility, and SWIPD-L, which uses a harmonic average. These mobility fluxes yield similar estimates for the coercivity condition and accuracy as previous schemes. Numerical experiments illustrate the performance of the proposed schemes, demonstrating that they achieve optimal convergence rates and that the harmonic mobility flux provides potentially significant improvements in stability and accuracy compared to the arithmetic mean. An $hp$-adaptive formulation has also been presented, showing significant reductions in complexity and degrees of freedom compared to the $h$-adaptive formulation while maintaining similar accuracy and physical fidelity. Future work includes extending the theoretical analysis to provide rigorous structure-preserving proofs for the SWIPD-L scheme in the discrete setting, exploiting the harmonic average to extend Theorem 1.1. Additionally, more challenging test cases, such as the rising bubble benchmark, will further validate the performance of the proposed schemes for coupled CHNS systems and assess adaptivity effects under more demanding conditions.